OpenStudy (anonymous):

(n-1)!-(n)! is equal to what and how ??

myininaya (myininaya):

\[(n-1)!-n(n-1)!=(n-1)!(1-n)\]

OpenStudy (anonymous):

the answer is nxn! how come ??

myininaya (myininaya):

lets test what you said n*n!=(n-1)!-n! ? let;s see if it works for n=2 2*2!=(2-1)!-2! 4=1-2 <- not true now lets test what i said (n-1)!(1-n)=(n-1)!-n! ? let's try n=2 (2-1)!(1-2)=(2-1)!-2! 1(-1)=1-2 -1=-1 <- this is true

myininaya (myininaya):

so n* n! is total not right

myininaya (myininaya):

totally*